given the production Q=L^1/2 K^1/2 price per unit of K and L is birr 2 and birr 4 respectively and total cost is birr 80 a) determine the maximum output subject to the cost constraint b)caculate the MRTSl,k at equilibrium point


1
Expert's answer
2021-12-10T11:25:21-0500

The optimal L and K will be at the point where

"MRTS=\dfrac{MPL}{MPK}=\dfrac{w}{r}"


Therefore

"MPL=\dfrac{1}{2}L^{-1/2}K^{1/2}\\[0.4cm]
MPK=\dfrac{1}{2}L^{1/2}K^{-1/2}\\[0.4cm]"

If "w=4 \text{ and } r=2" , then


"\dfrac{\frac{1}{2}L^{-1/2}K^{1/2}}{\frac{1}{2}L^{1/2}K^{-1/2}}=\dfrac{4}{2}\\[0.4cm]

\dfrac{K}{L}=2\\[0.4cm]
K=2L"


The total cost is equal to

"80=4L+2K"


Therefore

"80=4L+2(2L)\\[0.4cm]
80=4L+4L\\[0.3cm]
80=8L\\[0.3cm]
L^*=\boxed{10}"

The optimal capital is equal to

"K=2\times 10\\[0.3cm]
K^*=\boxed{20}"


At the optimal L and K, the MRTS is equal to


"MRTS=\dfrac{10}{5}\\[0.3cm]
MRTS=\boxed{2}"